# SofT-GRPO: Surpassing Discrete-Token LLM Reinforcement Learning via Gumbel-Reparameterized Soft-Thinking Policy Optimization

### v1

@@ Appendix D Supplementary of Experiments
D.1 Supplementary of Token Efficiency
As shown in Section 4.5, besides the performance, the token efficiency of LLMs is also an important metric. In this section, we compare the token efficiency of baselines and SofT-GRPO in Table 6. Compared to No-Finetune variants, SofT-GRPO can demonstrate a clear refinement in both the token efficiency across all queries and the token efficiency across correct queries. Compared to discrete-token GRPO, SofT-GRPO will not cause severe token improvement. Specifically, we observe a severe reduction in the thinking length of the LLaMA-3.2-3B-Instruct model. As shown in Figure 4, unlike GRPO, SofT-GRPO maintains and even enhances the token efficiency compared to the base LLM when training progresses, demonstrating the effectiveness of SofT-GRPO in reducing computational consumption.
@@ D.1 Supplementary of Token Efficiency
D.2 Experiments under Different Temperatures
In Section 4, we adopt the setting of $\tau=0.6$ for discrete-token CoT. To investigate whether other temperature settings will break our observations in our main experiments, similar to yue2025does, we try temperatures from $\tau=0.6$, $\tau=0.8$, $\tau=1.0$, $\tau=1.2$, and $\tau=1.4$.
@@ D.2 Experiments under Different Temperatures
D.3 Supplementary of Ablation on Hyper-Parameters
@@
(b) Training Proximal Policy Optimization (PPO) (schulman2017proximal) KL Divergence curve between $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta}$
Figure 6: KL Divergence curves in SofT-GRPO.
In this subsection, we have a further investigation into the hyperparameter settings. As briefly shown in Section 5.2, adopting a higher top-p or $\tau_{g}$ will cause collapses in training. We attribute these collapses to the case that some soft-thinking tokens may become incomprehensible for LLMs.
@@
Recently, qi2025defeating provides excellent insight into the collapse situation in the RLVR fine-tuning. When collapse is caused by a precision issue, they observe a super high KL divergence between $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta}$ (Refer to Figure 3 in (qi2025defeating), $10^{-3}$ even higher). However, in the variants shown in Figure 6(b), we find that their KL divergence in PPO policies is less than $10^{-5}$, indicating that the variants of SofT-GRPO are less likely to meet precision issues.
Appendix E Baselines & Datasets & Licenses
@@
E.1 Baselines
We mainly include No-Finetune base LLMs, Discrete-Token GRPO, and the method in butt2025soft (noted Soft Token) as baselines.
@@ E.1 Baselines
We utilize the latest verl (sheng2024hybridflow) https://github.com/volcengine/verl/tree/main and vLLM rollout to implement discrete-token GRPO with default parameters.
butt2025soft Soft Token
Due to the requirement of passing d-dimensional inputs $\hat{\boldsymbol{s}}$ between the rollout workers and the verl policy optimization workers, implementing this algorithm requires a high amount of communication between the rollout workers and the policy update workers. So, we report the results in (butt2025soft) for comparison in Table 2.
E.2 Datasets
This paper covers five in-domain numerical reasoning datasets (i.e., AIME2024, AIME2025, AMC23, MATH-500, and GSM8K (cobbe2021training)), one out-of-domain scientific reasoning dataset GPQA-Diamond (rein2024gpqa), and two out-of-domain code reasoning datasets (i.e., HumanEval and MBPP). These datasets are provided in https://github.com/eric-ai-lab/Soft-Thinking. Specifically, due to their answer can be correct with various equivalent forms, we exclude the MATH-500 benchmark for Table 4.
E.3 Licenses

### v2

@@ Appendix E Supplementary of Experiments
E.1 Supplementary of Token Efficiency
As shown in Section 4.6, besides the performance, the token efficiency of LLMs is also an important metric. In this section, we compare the token efficiency of baselines and SofT-GRPO in Table 9. Compared to No-Finetune variants, SofT-GRPO can demonstrate a clear refinement in both the token efficiency across all queries and the token efficiency across correct queries. Compared to the discrete-token GRPO, SofT-GRPO will not cause severe token improvement. Specifically, we observe a severe reduction in the thinking length of the LLaMA-3.2-3B-Instruct model. As shown in Figure 4, unlike GRPO, SofT-GRPO maintains and even enhances the token efficiency compared to the base LLM when training progresses, demonstrating the effectiveness of SofT-GRPO in reducing computational consumption.
@@ E.1 Supplementary of Token Efficiency
E.2 Experiments under Different Temperatures
In Section 4, we adopt the setting of $\tau=0.6$ for discrete-token CoT. To investigate whether other temperature settings will break our observations in our main experiments, similar to Yue et al. (2025), we try temperatures from the collection of $\tau=0.6$, $\tau=0.8$, $\tau=1.0$, $\tau=1.2$, and $\tau=1.4$.
@@ E.2 Experiments under Different Temperatures
E.3 Supplementary Ablation on Hyper-Parameters
@@
(b) Training Proximal Policy Optimization (PPO) (Schulman et al., 2017) KL Divergence curve between $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta}$
Figure 6: KL Divergence curves in SofT-GRPO over DeepSeek-R1-Distill-Qwen-1.5B Base LLM.
In this subsection, we further investigate the hyperparameter settings. As briefly shown in Section 5.2, adopting a higher top-p or $\tau_{g}$ will cause collapses in training. We attribute these collapses to the case that some soft-thinking tokens may become incomprehensible for LLMs.
@@
Recently, Liu et al. (2025a); Qi et al. (2025b) provides excellent insight into the collapse situation in the RLVR fine-tuning. When collapse is caused by a precision issue, they observe a very high KL divergence between $\pi_{\theta_{\text{old}}}$ and $\pi_{\theta}$ (Refer to Figure 3 in (Qi et al., 2025b), $10^{-3}$ even higher). However, in the variants shown in Figure 6(b), we find that their KL divergence in PPO policies is less than $10^{-5}$, indicating that the variants of SofT-GRPO are less likely to have precision issues.
E.4 Pass@K up to K=1024
Table 1 and Figure 5 show the Pass@K for K at most 64. In Yue et al. (2025), they conduct experiments for K up to 1024. We adopt this setting. Confined by limited computational resources, we implement the proposed SofT-GRPO over DeepSeek-R1-Distill-Qwen-1.5B for 1024 runs, as well as the GRPO and No-Finetune baselines with the discrete-token CoT reasoning pattern.
As shown in Figure 7, we calculate Pass@16 to Pass@1024 on AIME2024 and AMC 23. As a difference from the result in Table 1, we obtain Pass@K for each K runs instead of the first K runs. For example, when calculating Pass@32, we divide the 1024 runs into 32 groups, calculating their Pass@K value and averaging. In Table 1 instead, we calculate the Pass@K at first 32 runs.
Figure 7 obtains the same result pattern compared to the results in (Yue et al., 2025), where GRPO will outperform the No-finetune version when K is small, but will be left behind when K increases. SofT-GRPO can consistently outperform the GRPO and no-finetune counterparts on all K (from 16 to 1024).
(a) Pass@16 to Pass@1024 on AIME2024
(b) Pass@16 to Pass@1024 on AMC23
Figure 7: Pass@16 to Pass@1024 on AIME2024 and AMC 23. The fluctuation in Figure (a) occurs because AIME only has 30 instances, and SofT-GRPO correctly solved three out of four (1024/256) problems when K=256. In Figure (b), all algorithms ultimately achieved complete correctness. Consistent with this, we checked CoT to prove that the correct results did not come from random guessing.
E.5 P-values for Significance on Pass@K
Besides the results in Table 1 and Figure 5, in this section, we use results from 1024 runs to demonstrate the significance test for Pass@16 and Pass@32 (dividing 64 groups for Pass@16 from the 1024 runs and 32 groups of Pass@32). We calculated the t-test for the one-sided hypothesis test (See Figure 7, the average result of SofT-GRPO outperforms the GRPO and No-Finetune baselines), and taking 0.05 as the threshold, the results in Figure 8 and 9 show that in Pass@16 with 64 runs, and Pass@32 with 32 runs:
  • SofT-GRPO can significantly outperform GRPO on AIME2024 with Pass@16 (0.0067), AMC23 with Pass@32 (0.022), with quite small values (0.13) on both AIME2024 with Pass@32 and AMC23 with Pass@16.
  • SofT-GRPO can get quite small p-values (0.081 and 0.054) on AMC23 with Pass@16 and Pass@32.
(a) P-scores of Pass@16 over SofT-GRPO, GRPO, and No-finetune on AIME2024
(b) P-scores of Pass@32 over SofT-GRPO, GRPO, and No-finetune on AIME2024
Figure 8: Significance test on AIME2024.
(a) P-scores of Pass@16 over SofT-GRPO, GRPO, and No-finetune on AMC23
(b) P-scores of Pass@32 over SofT-GRPO, GRPO, and No-finetune on AMC23
Figure 9: Significance test on AMC23.
Appendix F Discussion & Analysis on Results
In this section, we will provide a detailed analysis of the improvement in Pass@K. Yue et al. (2025) first validates that RLVR techniques in discrete-token GRPO will not incentivize reasoning capacity in LLMs beyond the base model, demonstrating equal or even inferior Pass@K values when k improves. So, Recent studies (Walder and Karkhanis, 2025; Mahdavi et al., 2025; Peng et al., 2025) try to develop RL methods that optimize the Pass@K accuracy directly. Among them, Walder and Karkhanis (2025); Chow et al. (2024); Mahdavi et al. (2025); Hu et al. (2026) focus on reallocating the reward within the group, encouraging rewards of negative discrete-token CoTs in groups with high Pass@K value. Peng et al. (2025) try to improve the model’s Pass@K performance by avoiding the over-concentration on the top-1 selection within the token set.
In SimKO, at high-entropy (forking) tokens, positive (correct) samples are rewarded not only on the sampled token but spread across the top-K most probable tokens, promoting output diversity and better Pass@K. For negative (incorrect) samples, SimKO applies stronger penalties to the top-1 token and lighter penalties to others, which prevents the output probabilities from collapsing onto a single choice. This entropy-aware and asymmetric strategy encourages exploration, outperforming traditional RL methods that only focus on the sampled token.
In the proposed SofT-GRPO, we have the policy update in the soft-thinking part ($t$-th token) as:
$\nabla_{\theta}L(\theta)$ $=\nabla_{\theta}\left(\sum_{i=1}^{|\mathcal{T}|}\left[-g^{\prime}_{i}+\log p_{i}-\exp(-g^{\prime}_{i}+\log p_{i})\right]-\left[-\epsilon_{i}-\exp(-\epsilon_{i})\right]\right)$
$=\sum_{i=1}^{|\mathcal{T}|}\nabla\theta\left(\log p_{i}-\exp(-g^{\prime}_{i}+\log p_{i})\right)$
$=\sum_{i=1}^{|\mathcal{T}|}\left(\nabla_{\theta}\log p_{i}-\exp(-g^{\prime}_{i}+\log p_{i})\nabla_{\theta}\log p_{i}\right)$
$=\sum_{i=1}^{|\mathcal{T}|}\left(1-\exp(-g^{\prime}_{i}+\log p_{i})\right)\nabla_{\theta}\log p_{i}$
$=\sum_{i=1}^{|\mathcal{T}|}\left(1-p_{i}\cdot\exp(-g^{\prime}_{i})\right)\nabla_{\theta}\log p_{i}.$
This means that instead of updating the policy only for the selected token, in SofT-GRPO, it inherently highlights: (1). Higher importance component ($g^{\prime}_{i}$) in the soft tokens, (2). lower prob ($p_{i}$). For example, for positive instances, SofT-GRPO inherently improves the probability of the main component in each soft-tokens and the low-probability components. For Negative instances, SofT-GRPO inherently gives penalties to the probability of the main component in each soft-tokens and the low-probability components.
Figure 10 verifies this conclusion: the entropy of token probability distributions will converge to nearly 0 with discrete-token GRPO, but keep stable and even improve with SofT-GRPO.
Figure 10: Entropy curve on DeepSeek-R1-Distill-Qwen-7B during training.
Appendix G Baselines & Datasets & Licenses
@@
G.1 Baselines
We mainly include No-Finetune base LLMs, Discrete-Token GRPO, and the method in Butt et al. (2025) (noted Soft Token) as baselines.
@@ G.1 Baselines
We utilize the latest verl (Sheng et al., 2024) https://github.com/volcengine/verl/tree/main and vLLM rollout to implement discrete-token GRPO with default parameters.
Butt et al. (2025) Soft Token
Due to the requirement of passing d-dimensional inputs $\hat{\boldsymbol{s}}$ between the rollout workers and the verl policy optimization workers, implementing this algorithm requires a high amount of communication between the rollout workers and the policy update workers. So, we only report the results in (Butt et al., 2025) for comparison in Table 4.
G.2 Datasets
This paper covers five in-domain numerical reasoning datasets (i.e., AIME2024, AIME2025, AMC23, MATH-500, and GSM8K (Cobbe et al., 2021)), one out-of-domain scientific reasoning dataset GPQA-Diamond (Rein et al., 2024), and two out-of-domain code reasoning datasets (i.e., HumanEval and MBPP). These datasets are provided in https://github.com/eric-ai-lab/Soft-Thinking.
G.3 Inference Framework
The inference framework of our implementation is built on SGLang (Zheng et al., 2024), maximizing efficiency via continuous batching, RadixAttention for KV cache reuse, and compressed finite state machines for faster structured output decoding. In the standard flow, the scheduler flattens discrete token inputs into continuous tensors for GPU execution; generated tokens are returned to update the Radix tree state for subsequent steps.
SGLang’s optional overlap scheduler pipelines execution by issuing the next batch ahead via negative integer slot addresses. Compared to Zhang et al. (2025b), we enabled this for Soft-Thinking by storing generated soft probabilities and indices in VRAM-allocated future maps. A custom kernel resolves the slot addresses to fetch these states just-in-time. This design should be faster by eliminating PCIe bottlenecks, allowing non-blocking scheduling with data kept entirely on the GPU.
G.4 Licenses
@@
Appendix H Visualization
In this section, in Figure 11 and Figure 12, we provide visualization examples for the outcome of SofT-GRPO on two examples (we show $y_{i}$ numbers that are larger than 0.01 in soft-tokens). Our observations are as follows:
  • Compared to Figure 4 in Zhang et al. (2025b), SofT-GRPO preserves the multiply-path thinking pattern, thus providing better soft-thinking reasoning performances.
  • As discussed in Wu et al. (2025a), SofT-GRPO outcomes can observe possible latent search trees. See the 288th soft-token in Figure 11, it blends two discrete tokens with contrary meanings in its top-2 (No and Yeah, probably with totally different embeddings), representing a more abstract concept than any language token. Moreover, to check whether this pattern has already existed in No-Finetune LLM, we searched their output on GSM8K (a total of 1319 instances), where we could not find this probability distribution. So, we can recognize that this probability distribution is highly possible to emerge after SofT-GRPO.
  • Similarly to Figure 4 in Zhang et al. (2025b), SofT-GRPO can preserve interpretability; the contents in the red box (Okay, so I need to…) can be a good explanation for the path of soft-thinking reasoning.
Figure 11: Visualization: DeepSeek-R1-Distill-Qwen-1.5B + SofT-GRPO on the 4-th instance of GSM8K.
Figure 12: Visualization: DeepSeek-R1-Distill-Qwen-1.5B + SofT-GRPO on the 4-th instance of MATH-500.